OpenStudy (anonymous):

Twelve students in a class average 70% on a certain test. Eighteen others average 80%. What is the overall average of the thirty students as a percent?

OpenStudy (anonymous):

if the total score of the 12 student class is \(x\) then \(\frac{x}{12}=70\) and \(x=70\times 12=840\) similarly if the total test score of the 18 student class is \(y\) then \(\frac{y}{18}=80\) so \(y=18\times 80=1440\) the total combined is \(1440+840=2280\) so the average over the 30 students is \(\frac{2280}{30}=76\)

OpenStudy (anonymous):

thanks man !

OpenStudy (anonymous):

yw